### v1

# Fixing Code Generation Errors for Large Language Models

Abstract
Code generation leverages artificial intelligence technologies, particularly Large Language Models (LLMs), to automatically produce source code, enhancing software development efficiency and reducing repetitive tasks. However, the LLMs’ generated code often fails to pass test cases and requires substantial human effort to fix errors. Previous studies focused on better prompts or improving LLMs’ capability but ignored why LLMs failed. In this paper, we first reproduced 14 LLMs, including GPT-3.5-turbo and 13 open-source LLMs, on the HumanEval dataset. We extracted 12,837 code generation errors and conducted an in-depth analysis of their causes, which led to the identification of 19 distinct error causes. Our empirical analysis indicated that three of these causes can be directly fixed. Consequently, we proposed a fixing method called LlmFix, which addresses these three types of errors through a three-step process: filtering code for indentation correction, truncating redundant generated code, and importing missing modules. Experimental results demonstrate that LlmFix can fix these three types of errors, significantly improving the performance of 14 LLMs on HumanEval and MBPP datasets with average increases of 9.5% and 5.4%, respectively.
Index Terms:
Large Language Model, Code Generation, Automatic Fixing
I Introduction
Code generation refers to the process of converting natural language descriptions of software development tasks into automatically generated code snippets that effectively solve these specific tasks [1, 2]. In real-world applications, developers expect the generated code to not only compile successfully but also perform the intended tasks accurately, in alignment with specified requirements or problem statements. Effective code generation can significantly reduce developers’ time and effort in coding, enabling them to concentrate on more complex and innovative aspects of software engineering [3, 4].
Large Language Models (LLMs) based on Transformer architecture have become powerful tools for various natural language processing (NLP) tasks, including sentiment analysis [5, 6] and language translation [7]. LLMs pre-trained on extensive codebases and fine-tuned specifically for code generation have also demonstrated promising performance in generating code [2, 8]. Furthermore, Zheng et al. [9] conducted a comprehensive survey of existing open-source code generation LLMs, reporting their performance on the widely used HumanEval dataset.
Despite advancements in LLMs for code generation, their output often contains errors, prompting researchers to explore various methods for improvement. Liu et al. [10] evaluated ChatGPT’s capabilities for two code generation tasks, including text-to-code and code-to-code generation. They designed prompts by leveraging the chain-of-thought strategy with multi-step optimizations. Their findings demonstrated that carefully crafted prompts for ChatGPT can significantly enhance the performance of code generation. White et al. [11] explored how to optimize interactions with LLMs by designing efficient prompt patterns, highlighting the importance of crafting effective prompts. They also provided concrete implementation examples and demonstrated the practical effects of these patterns. Chen et al. [12] formalized an algorithm for learning from natural language feedback at training time, which they call imitation learning from language feedback.
While previous studies improve code generation performance by refining prompts or enhancing the model itself, they do not investigate the root causes of test case failures in model-generated code, nor do they explore possible methods to fix these errors directly after the generation. However, understanding and fixing these errors is crucial to achieving efficient code generation.
In this paper, we first conduct an empirical study to investigate the performance of LLMs in code generation, and identify the types of errors they generate. Specifically, we select 14 LLMs and, referencing the experimental results of Zheng et al. [9] on the HumanEval dataset [13], accurately reproduce their code generation performance on the HumanEval dataset, followed by a series of in-depth empirical studies. Our findings reveal that LLM performance generally correlates positively with parameter scale, while inference time shows no significant association with parameter size. Then, to investigate the reasons behind LLMs failing in code generation tasks, we add “try-except” code in each generated code and collect 12,837 errors related to 14 types of exceptions (e.g., AssertionError, NameError, etc.). According to in-depth manual analysis, we ultimately identify 19 specific causes leading to these errors. Our research reveals that 16 causes (e.g., Inappropriate Argument, Unbalanced Delimiters, etc.) belong to the incapability of LLMs, and only three causes were amenable to direct automated fixing including: 1) Inconsistent Indentation, the code fails to pass the compilation due to wrong indentation; 2) Function Overflow, the completed function is appended many redundant and partial code as LLM does not stop in time; 3) Missing Import, the generated code invokes modules without import.
Based on these empirical research findings, we design LlmFix, a method specifically tailored to automatically fix errors caused by the above three causes, which includes the following three corresponding operations: 1) Code Filtering, searching for inconsistent indentation and replacing it with uniform tab-based indentation. 2) Code Truncation, discarding redundant code located after the function completion. 3) Importing Missing Modules, automatically adding corresponding import statements at the beginning of the code when detecting the use of a module that has not been imported. We evaluate our LlmFix on two datasets (i.e., HumanEval and MBPP). Experimental results demonstrate that LlmFix can fix these three types of errors which are amenable for automated fixing and improve the performance of the 14 LLMs by an average of 9.5% and 5.4% on two code evaluation benchmarks, respectively.
In summary, the major contributions of this study are:
  • $\bullet$
    Comprehensively investigating the code generation performance of 14 LLMs, analyzing the relationship between model performance, inference time, and parameter size.
  • $\bullet$
    Systematically categorizing and analyzing 12,837 code generation errors to reveal error distributions and reasons.
  • $\bullet$
    Proposing LlmFix for automatically fixing code generation errors and demonstrating its effectiveness of 14 LLMs on HumanEval and MBPP datasets.
  • $\bullet$
    Sharing our replication package [14].
This paper is structured as follows. Section II introduces the related work on code generation. Section III provides the empirical study on LLM replication and error analysis. Section IV presents our error fixing method LlmFix, which is evaluated in Section V. Section VI provides implications for developers and researchers and the threats to the validity of our study. Section VII summarizes this paper.
II Related Work
This section describes the existing code generation studies in subsection II-A. It then provides the relevant LLM for code generation in subsection II-B, leading to the goal of this study.
II-A Code Generation
The task of code generation has evolved significantly over time, and researchers have developed many solutions [15, 16, 17, 18, 19]. Early methods can quickly generate simple code structures based on predefined templates [20] or rule systems [21]. Syriani et al. [17] categorized template-based code generation approaches, identifying their prevalent characteristics and trends, emphasizing the technique’s significance in generating source code across diverse application areas. These simple methods are intuitive to implement and control. However, they have difficulties in implementing complex code and generating code patterns not explicitly defined in the rules [22].
As deep learning (DL) techniques evolved, researchers leveraged approaches such as long short-term memory (LSTM) [23], to comprehend code semantics. By learning from numerous code data, e.g., from the open-source community GitHub, these models can capture the structural and semantic patterns of programming languages, and generate more complex code [24]. Yin et al. [16] introduced a new neural architecture that incorporates a grammar model to explicitly integrate the target syntax as prior knowledge. Their experiments demonstrate that this method is an effective strategy for scaling up to the generation of complex programs from descriptions in natural language. Generally, the DL-based models can produce syntactically correct code, but the generated code was commonly not fully aligned with the given task description [25].
In recent years, the emergence of large language models (LLMs) like GPT [26] and Codex [13] have introduced a new phase in code generation. These models are based on the Transformer architecture [27] and pre-trained on large-scale code datasets to learn the deep semantics and syntactic rules of programming languages. LLMs can better understand developers’ programming requirements in natural language and generate more satisfactory code compared to existing DL-based models [10]. Alawida et al. [28] pointed out that LLMs like ChatGPT strongly influenced the software industries in terms of their architecture, training data, evaluation metrics, etc. Moreover, Zheng et al. [9] reviewed state-of-the-art LLMs and compared their performance for code generation. Unfortunately, they did not uncover why these LLMs failed in code generation.
II-B LLM for Code Generation
With the rise of LLMs exemplified by ChatGPT, many researchers have conducted studies on LLMs-based code generation [29, 30, 31]. These studies encompass various aspects. Previously, many studies focused on building a high-quality evaluation dataset for code generation. To address this issue, Lu et al. [29] introduced CodeXGLUE, a benchmark with 14 datasets to foster DL-based research for program understanding and generation. However, this dataset used evaluation measures like BLEU [32], which cannot verify whether an LLM can generate functionally correct code. Li et al. [33] proposed an open-source large-scale code generation dataset TACO which includes rigorous data processing and unit test validation, although this dataset only focused on competition-level programming questions. Chen et al. [13] proposed HumanEval, an innovative collection of 164 real-world programming challenges in Python. The evaluation encompasses not only syntax but also logical correctness and problem-solving efficiency, requiring the generated code to pass many test cases. HumanEval has made this dataset a widely used benchmark for code generation evaluation [13].
Based on the above evaluation datasets, researchers conducted studies on the factors affecting LLMs’ performance in code generation [34, 35, 36]. Liu et al. [10] found that the performance is sensitive to the LLMs’ prompts, namely the programming requirement written in natural language as model input. By leveraging a chain-of-thought strategy, the optimized prompts can lead to better code generation performance. White et al. [11] focused on improving interactions with LLMs through efficient prompt patterns. They found that prompt patterns significantly enrich the capabilities that can be created in a conversational LLM and the prompt patterns are generalizable to many different domains. Chen et al. [12] introduced Imitation Learning from Language Feedback (ILF). They have shown that ILF can significantly improve the quality of a code generation model, even with just a small sample of human-written feedback and refinements. Diao et al. [37] proposed a new method, Active-Prompt, to adapt LLMs to different tasks with task-specific example prompts. Active-Prompt achieved a promising performance on eight widely used datasets for arithmetic reasoning, commonsense reasoning, and symbolic reasoning. Zhao et al. [38] proposed Logical Thoughts (LoT) to improve the zero-shot chain-of-thought reasoning ability of large language models. Their experiments demonstrate that the method leads to improved reasoning ability, particularly when applied to LLMs of large scale. In summary, existing studies mainly focus on prompt optimization and have not investigated why LLMs did not pass the test cases. To fill this gap, this study aims to analyze why LLMs failed and explore the way to directly fix the generated code.
III Empirical Study
III-A Research Questions
This empirical study aims to systematically analyze and compare the performance of various LLMs, and to categorize the errors in the generated results to guide potential fixing methods. To achieve this goal, we investigate the following RQs:
RQ1: Can we reproduce the performance of state-of-the-art LLMs? This RQ examines the reproducibility of results claimed in the literature on LLMs for code generation. By attempting to replicate the conditions and methodologies of previous studies, we aim to assess the validity of reported performance metrics. This RQ is crucial for establishing a baseline of expectations when it comes to the operational capabilities of LLMs.
RQ2: Why did these LLMs fail in code generation? Investigating the code generation errors of LLMs is essential for limitation understanding and further improvements. By running the generated code with test cases and analyzing the occurred errors, we aim to uncover the underlying factors that contribute to these errors.
RQ3: Which errors can be directly fixed? Improving the code generation performance is the final goal of this study. This RQ aims to analyze the possibility of automatically fixing some errors after the code generation without asking users to update prompts or sending regeneration requests.
III-B Dataset and Evaluation Measure
To conduct the empirical study, we utilize the widely used dataset HumanEval [13], meticulously crafted by OpenAI, for assessing various LLMs. The HumanEval dataset is an innovative collection of 164 real-world programming challenges in Python. Each generation task involves an average of 7.7 test cases, ensuring the functional correctness of the generated code. Most of the well-known LLMs such as CodeLlama [19], CodeGen [39], and WizardCoder [40] utilized HumanEval for evaluating code generation capabilities. As shown in Fig. 1, the HumanEval comprises five components: 1) task_id, a unique identifier for each task, allowing for easy reference and organization of the code generation tasks. 2) prompt, a textual statement that provides generation requirements for guiding LLMs. 3) canonical_solution, the standard solution for each task. 4) test, a testing function along with several test cases, designed to meticulously assess whether the generated code meets the expected requirement. 5) entry_point, the name of a function, which is intended to be the main entry point of the generated code function.
Figure 1: A schematic diagram of the first sample in HumanEval.
For this dataset, an LLM needs to generate the functionally correct code, which can pass all the test cases, according to a given prompt. To measure the accuracy of code generation, Chen et al. [13] provide the $Pass@k$ metric, the percentage of tasks solved by an LLM. A task is considered solved if any top-$k$ code generated by LLM can pass all the test cases. As Pass@k has high variance, we follow Chen et al. [13] and use the unbiased version: $Pass@k=\mathbb{E}\left[1-\binom{n-c}{k}/\binom{n}{k}\right]$, where $\mathbb{E}$ denotes the average performance across all tasks; $n$ is the total number of tasks; $c$ is the number of correctly solved tasks; $\binom{n}{k}$ is the number of combinations for choosing $k$ tasks out of $n$; $\binom{n-c}{k}$ is the number of combinations for choosing $k$ tasks from those one model cannot solve (i.e., $n-c$ tasks), representing the number of ways to fail. Note that this study measured the code generation performance with $Pass@1$, as developers commonly expected that LLMs could generate correct code in the first place. Besides, to mitigate the effects of randomness in LLMs, we follow [41] to test each LLM ten times and then report the average value as the final $Pass@1$ result.
III-C Can We Reproduce the Performance of State-of-the-Art LLMs? (RQ1)
III-C1 Motivation and Method
In recent years, researchers have developed many LLMs and published their test results on the HumanEval dataset [9]. However, achieving outcomes that are completely identical to those reported remains a challenging task [42]. Therefore, to thoroughly investigate whether existing LLMs are consistent with the reported performances as described in RQ1, we adopted several state-of-the-art LLMs for testing. One is the ChatGPT, a state-of-the-art closed-source LLM released by OpenAI [43]. Under zero-shot conditions, ChatGPT has achieved an accuracy rate of over 60% on the HumanEval dataset maintaining a leading position [44].
For open-source LLMs, Zheng et al. [9] reported on the performance of many LLMs on HumanEval. Based on this study, we selected four families of LLMs. Phind [45] performs best in the reported ranking, followed by WizardCoder [40] and CodeLlama [19]. Phi [46] is the best small-sized LLM and is competitive with other families. Besides, we included the CodeTulu family [47] in our study due to its promising performance as reported by Ivison et al. [47]. All open-source LLMs are re-implemented by using the LLM parameters stored in the HuggingFace platform and tested on a server with four NVIDIA RTX3090 GPUs in default settings.
Detailed information about each LLM family and relevant links are listed below. We ran each LLM ten times on HumanEval and compared the code generation performance from original reports [9, 47].
  • $\bullet$
    ChatGPT, is a powerful LLM released by OpenAI [48]. Researchers have demonstrated its high performance in code generation [49]. We chose the GPT-3.5-turbo version with 175B parameters by invoking official APIs [50] following Sun et al. [51].
  • $\bullet$
    Phind, was introduced by the Phind team, available in two versions [52]: Phind-CodeLlama-34B-v1 and Phind-CodeLlama-34B-v2. According to the team’s report [45], the latter version even outperforms OpenAI’s GPT series in code generation.
  • $\bullet$
    WizardCoder, was introduced by Microsoft [40], which empowers code generation LLMs with complex instruction fine-tuning. We re-implemented the versions of WizardCoder-Python in different parameter sizes (7B/13B/34B) [53].
  • $\bullet$
    CodeLlama, was introduced by Meta AI [19], which is a competitive code generation model and inspired many other varied versions such as CodeTulu [47] and WizardCoder [40]. We used the CodeLlama-python (7B/13B/34B) [54] as our baseline, which fine-tunes CodeLlama with more Python code and shows better performance in code generation.
  • $\bullet$
    Phi, was released by Microsoft [46], which is renowned for its smaller model size (1.3B) but demonstrated exceptional high performance in code generation. Later, Microsoft released new versions including Phi-1.5 Li et al. [55] and Phi-2 [56]. Thus, these three models were selected as our baselines including Phi-1 (1.3B), Phi-1.5 (1.3B), and Phi-2 (2.7B) [57].
  • $\bullet$
    CodeTulu, was proposed by AllenAI [58], whose coding capabilities matched or even exceeded the performance of GPT-3.5-turbo in several benchmark tests. We selected the best-performing and variably parameterized versions of CodeTulu-2 (7B/13B/34B) as our baselines.
III-C2 LLMs Reproduction and Evaluation
Fig. 2 shows the accuracy distribution of 14 LLMs over ten times tests, sorted by the average accuracy from lowest to highest. The best-performing model is the Phind-CodeLlama-V2-34B, with an average accuracy of 63.29%. Table I lists the specific numerical characteristics of these 14 LLMs from the ten tests, including minimum, maximum, mean, standard deviation, and results reported in related research papers [9]. From the study by Zheng et al. [9], we know that the 14 LLM accuracies range from 34.10% to 73.80%, while our replicated test results have an average accuracy ranging from 27.14% to 63.29%. Compared to reported data, our average test results are approximately reaching 84.07% of the reported results effectively. This difference is within an acceptable range, likely related to the numerical precision used in the experiments. Additionally, the standard deviation (STD) for all models is significantly lower than the mean (less than 10% of the mean), indicating low data variability.
Figure 2: The distribution of test results for 14 models, sorted by the mean values from lowest to highest, where the thin vertical lines represent the means; the thick short lines inside the boxes represent the medians; and the stars indicate the reported performance in the related papers [9].
TABLE I: Accuracy of ten tests of LLMs on HumanEval dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values, compared to the result reported in related papers [9], where the percentage value in parenthesis indicates the improvement ratio to the reported; ”-” indicates the related paper did not report the Pass@1 as required in our study but Pass@10.
Model Reported Minimum Maximum Mean STD
CodeLlama-Python-7B 38.40 31.10 (--19.01%) 35.37 (--07.89%) 32.69 (--14.87%) 1.39
CodeLlama-Python-13B 43.30 31.71 (--26.77%) 40.24 (--07.07%) 36.65 (--15.36%) 2.82
CodeLlama-Python-34B 53.70 39.02 (--27.34%) 46.34 (--13.71%) 43.72 (--18.58%) 2.17
WizardCoder-Python-V1.0-7B 55.50 40.05 (--27.75%) 43.90 (--20.90%) 42.23 (--23.91%) 1.05
WizardCoder-Python-V1.0-13B 64.00 48.17 (--24.73%) 56.10 (--12.34%) 52.13 (--18.54%) 2.48
WizardCoder-Python-V1.0-34B 73.20 57.32 (--21.72%) 62.80 (--14.21%) 59.09 (--19.29%) 1.58
Phind-CodeLlama-V2-34B 73.80 60.37 (--18.20%) 65.24 (--11.60%) 63.29 (--14.24%) 1.51
Phi-1-1.3B 50.60 43.29 (--14.45%) 46.95 (--07.21%) 45.30 (--10.47%) 1.04
Phi-1.5-1.3B 34.10 23.78 (--30.26%) 28.66 (--15.95%) 27.14 (--20.41%) 1.42
Phi-2-2.7B - 29.27 40.24 35.55 3.13
GPT-3.5-turbo 60.30 55.49 (--07.98%) 60.37 ($+$00.12%) 58.11 (--03.63%) 1.46
CodeTulu-2-7B - 31.71 37.20 33.97 1.55
CodeTulu-2-13B - 37.20 40.85 38.84 1.38
CodeTulu-2-34B - 44.51 51.22 48.66 2.30
To assess reproducibility, we calculated the Pearson correlation coefficient [59] between our mean values and the reported values. The correlation coefficient of about $0.989$ indicates almost perfect consistency between our results and the reported data. The corresponding P-value of $6.91\times 10^{-9}$ is highly significant statistically, demonstrating a very high level of replication. Phi-2 and the CodeTulu-2 family of LLMs did not report the usual Pass@1 values. Instead, the CodeTulu-2 family chose to report Pass@10 to highlight the superiority of their models. In our study, these models had an average accuracy of 39.26%, ranking ninth among all models.
Table II shows the numerical characteristics of inference times per sample for 13 types of open-source LLMs mentioned, over ten tests, including minimum, maximum, average, and standard deviation values. It is worth noting that GPT-3.5-turbo, being a closed-source model running on OpenAI’s servers, operates in a different testing environment from the other 13 open-source models. Consequently, we have excluded its inference time from this analysis. From the test results, the average inference times ranged from 16.84 seconds to 77.38 seconds, with the LLM Phi-1-1.3B having the shortest inference time and WizardCoder-Python-V1.0-34B having the longest. Additionally, the standard deviations (STD) of average inference times for most LLMs were significantly lower than their corresponding averages (less than 10% of the mean), indicating low variability in inference times across most models.
TABLE II: The average testing time used per sample in ten LLM tests conducted on the HumanEval dataset with a total of 164 samples, in terms of the minimum, maximum, mean, and standard deviation (STD).
Model Min Max Mean STD
CodeLlama-Python-7B 25.18s 29.26s 27.33s 1.10
CodeLlama-Python-13B 31.40s 48.73s 41.42s 4.46
CodeLlama-Python-34B 65.29s 82.66s 76.49s 6.02
WizardCoder-Python-V1.0-7B 21.98s 25.45s 24.03s 1.12
WizardCoder-Python-V1.0-13B 17.16s 19.93s 18.12s 0.87
WizardCoder-Python-V1.0-34B 72.94s 82.53s 77.38s 2.66
Phind-CodeLlama-V2-34B 24.73s 29.43s 26.91s 1.16
Phi-1-1.3B 16.33s 17.28s 16.84s 0.25
Phi-1.5-1.3B 16.27s 17.97s 17.14s 0.55
Phi-2-2.7B 21.39s 22.72s 22.18s 0.36
CodeTulu-2-7B 33.27s 34.49s 33.96s 0.35
CodeTulu-2-13B 38.40s 56.50s 47.23s 6.23
CodeTulu-2-34B 22.90s 30.10s 25.37s 2.18
III-C3 Impact Analysis
Fig. 3 presents a linear regression analysis of the impact of parameter size on both model performance and average inference times. Through this analysis, the figure demonstrates two key relationships. On the model performance, the regression equation $y=0.5772x+34.083$ with a goodness of fit $R^{2}=0.5379$ reveals that, generally, the larger the model’s scale, the better its performance. The moderate goodness of fit value suggests that this statistical result has a certain level of significance. The data analysis from Table I similarly shows that this trend applies to most LLMs. To further analyze the performance of LLMs with different parameter sizes, we presented a bar graph, Fig. 4, featuring four groups of LLMs. Among LLMs with more than 20B parameters, the top three performers—Phind-CodeLlama-V2-34B, WizardCoder-Python-V1.0-34B, and GPT-3.5-turbo—demonstrated comparable performance. For LLMs with parameters between 10B and 20B, the best performer was WizardCoder-Python-V1.0-13B. In the range of 5B to 10B parameters, the best performer was WizardCoder-Python-V1.0-7B. For LLMs with less than 5B parameters, the best was Phi-1-1.3B. We observed that some LLMs, despite their smaller scale, achieved or even surpassed the accuracy of larger-scale LLMs. For instance, Phi-1-1.3B, though small in scale, achieved an accuracy of 45.3%, exceeding that of the 34B Codellama-Python model, which had an accuracy of 43.72%. This demonstrates that the size of the model’s parameters is not the sole determinant of performance.
Figure 3: Scatter plot and linear fit curve of accuracy versus inference time for 13 LLMs.
(a) Accuracy of models with less than 5B parameters.
(b) Accuracy of models with parameters between 5B and 10B.
(c) Accuracy of models with parameters between 10B and 20B.
(d) Accuracy of models with more than 20B parameters.
Figure 4: Bar chart of accuracy for four groups of models
On the average inference times, the regression function $y=0.9496x+20.249$ with a goodness of fit $R^{2}=0.3806$ is displayed. Compared to the fitting for accuracy, this $R^{2}$ value is significantly lower, suggesting substantial random fluctuations in the data and weak explanatory power of the regression function for inference times. This is confirmed by the data in Table II. For example, Phind-CodeLlama-V2-34B and CodeTulu-2-34B average about 25 to 26 seconds, while the smaller-scale CodeLlama-Python-13B requires 41.42 seconds. This is speculated to be related to the length of content generated by the models. Observing the file sizes of model outputs, we found that the former models’ output files are approximately 2MB, compared to 3.3MB for the latter. This indicates that the former models can solve the same problem with shorter code. Consequently, the time required to generate this shorter content is also shorter, resulting in a shorter overall inference time.
{mdframed}
[nobreak=true] Answer to RQ1: All our reproduced LLMs exhibit close performance as the original reports. The accuracy of LLMs is positively correlated with the size of their parameters in general except for two small-sized LLMs: Phi-1-1.3B and WizardCoder-Python-V1.0-7B. The average inference time of LLMs is not correlated with LLM scale, large LLMs such as Phind-CodeLlama-V2-34B and CodeTulu-2-34B demonstrate significantly shorter inference times than smaller-scale LLMs.
III-D Why did These LLMs Fail in Code Generation? (RQ2)
III-D1 Motivation and Method
To investigate why these LLMs failed in code generation, as described in RQ2, we extracted the error information from failed test cases and developed a systematic process for error analysis. The HumanEval dataset consists of 164 code generation tasks, each accompanied by an average of 7.7 corresponding test cases [13]. Our objective is to take the generated code and test cases as input and return the test results, assessing whether the code passed or failed. Utilizing the model and dataset interfaces provided by HuggingFace, we obtained the code generated by the model for each sample in the dataset, along with the corresponding test cases, all presented in string format. These strings can be parsed and executed as a series of Python statements using Python’s built-in exec() function, thus enabling automated testing. If the generated code passes all test cases, the test continues running; otherwise, the generated code fails one or more test cases, reports an exception, and terminates.
To facilitate batch retrieval of test results and reported exception messages for in-depth error analysis, we employed try-except statements, placing the execution of the generated code and test cases within a try block to catch and record any exceptions. This approach allowed to capture of detailed information about the errors, including the exception type and specific exception information. The extracted information was structured into a specific data format to facilitate further analysis. As shown in Fig. 5, it consists of three main parts: 1) Task Identification: the current sample’s number and an indication of whether the model successfully generated code. 2) Code Implementation: the code content generated by the model, based on the prompt part of the dataset, which is a completion of the original prompt. 3) Test Result Information: a detailed explanation of the test results for the code generated by the model, including whether the test passed or failed. If the test fails, it also provides the error type and specific error information that caused the error.
Figure 5: After testing each sample in the dataset, the output file displays the structure of the samples, including an example of a sample that passed the test and another that did not pass the test.
III-D2 Error Classification and Causes Analysis
Through ten rounds of testing on 14 different LLMs using the HumanEval dataset, we collected a total of 12,837 errors. Based on the Test Result Information shown in Fig. 5, we categorized these errors into 14 Python exception types. Fig. 6 presents a three-dimensional bar chart illustrating the frequency of occurrence for each of these 14 error types across all LLMs over the ten test rounds. Moreover, considering that the same type of error can be triggered by various causes, we unanimously agreed that it is necessary not only to categorize errors but also to analyze the code containing errors, thereby identifying one or more causes leading to each type of error. The above research results are summarized in Table III. The following is a detailed explanation of each error type:
Figure 6: Distribution of total errors across 14 LLMs in ten test runs.
TABLE III: Error types and causes summary: categories, frequencies, causes, and code examples. Note that examples are provided in Appendix.
Error Type #Errors Error Cause One-line Description Example
AssertionError 8171(63.64%) Test Case Failure Model-generated function fails pre-set test cases listing A
Empty Function Model creates an empty function or uses pass statement listing A, A
NameError 2916(22.71%) Misremembered Name Model defines a function but calls it with incorrect name listing A
Missing Content Model doesn’t generate any content, leading to missing entry_point - [^1] The ”Missing content” error occurs when the language model fails to generate any code at all. In such cases, there is no code snippet to display as an example. This error is particularly challenging to illustrate but is included due to its significance in understanding model behavior.
Missing Import Model doesn’t import required modules or functions listing A
SyntaxError 739(5.76%) Unbalanced Delimiters Imbalance in quotes or parentheses in generated code listing A
Function Overflow Excessive function generation hits limit, causing incomplete output and SyntaxError. listing A
ValueError 337(2.63%) Empty Sequence Function fails to handle empty input, causing ValueError listing A
Intentional Raise Model includes raise statements for program robustness listing A
Inappropriate Argument Function receives correct type but inappropriate value listing A
IndexError 291(2.27%) Out of Bounds Attempting to access an index outside sequence range listing A
TypeError 220(1.71%) Incompatible Operation Applying operation to object of inappropriate type listing A
AttributeError 58(0.45%) Non-existent Attribute Accessing a non-existent attribute of an object listing A
TimeoutError 50(0.39%) Execution Timeout Code execution exceeds set time limit listing A
IndentationError 32(0.25%) Inconsistent Indentation Indentation levels in same code block are inconsistent listing LABEL:code_inden1
ModuleNotFoundError 11(0.09%) Missing Module Importing a non-standard library module that’s not installed listing LABEL:code_module
KeyError 7(0.05%) Non-existent Key Attempting to access a non-existent key in a dictionary listing LABEL:code_key
UnboundLocalError 2(0.02%) Unassigned Variable Referencing a local variable before assignment listing LABEL:code_Unbound
RecursionError 2(0.02%) Infinite Recursion Recursive function lacks proper termination condition listing LABEL:code_recur1
NotImplementedError 1(0.01%) Intentional Raise Model includes raise statements for program robustness listing LABEL:code_not
AssertionError. This error occurs when an assert() statement fails. It can be attributed to two specific causes. The first is “Test case failure”, which occurs when the model-generated function fails to pass the test cases in the HumanEval dataset (as illustrated in Listing A). The second is “Empty function”, which results from the model creating an empty function or using only a pass statement (Listings A and A).
NameError. This error occurs when a local or global name is not found. It can be attributed to three specific causes. “Misremembered name” occurs when the model defines a function but calls it with an incorrect name (Listing A). “Missing content” happens when the model doesn’t generate any content, leading to a missing entry_point. Due to the nature of this error, there is no code to display. “Missing import” is caused by the model not importing required modules or functions (Listing A).
SyntaxError. This error occurs when the parser encounters a syntax error. It can be attributed to two specific causes. “Unbalanced delimiters” occur due to an imbalance in quotes or parentheses in the generated code (Listing A). “Function Overflow” happens when the excessive function generation hits limit, causing incomplete output and SyntaxError (Listing A).
ValueError. This error occurs when a function receives an argument of the correct type but an inappropriate value. It can be attributed to three specific causes. “Empty sequence” occurs when a function fails to handle empty input (Listing A). “Intentional raise” happens when the model includes raising statements for program robustness (Listing A). “Inappropriate argument” is when a function receives a correct type but inappropriate value (Listing A).
Due to the lower frequency of occurrence for the following error, we had fewer samples to analyze, resulting in only one specific cause for each. IndexError occurs when attempting to access an index outside the sequence range (Listing A). TypeError happens when applying an operation to an object of an inappropriate type (Listing A). AttributeError is raised when accessing a non-existent attribute of an object (Listing A). TimeoutError occurs when code execution exceeds the set time limit (Listing A). IndentationError happens when indentation levels in the same code block are inconsistent (Listing LABEL:code_inden1). ModuleNotFoundError is raised when importing a non-standard library module that’s not installed (Listing LABEL:code_module). KeyError occurs when attempting to access a non-existent key in a dictionary (Listing LABEL:code_key). UnboundLocalError happens when referencing a local variable before assignment (Listing LABEL:code_Unbound). RecursionError is raised when a recursive function lacks a proper termination condition (Listing LABEL:code_recur1). Finally, NotImplementedError occurs when the model includes raise statements for program robustness (Listing LABEL:code_not), which is consistent with one of the causes leading to ValueError.
III-D3 Distribution of Error Types
Based on Fig. 4 and Table III, We have also found that the 14 types of errors can be categorized into three groups based on their frequency of occurrence, as illustrated in Fig. 7:
Figure 7: Taxonomy of errors introduced while generating code using LLM.
From the perspective of errors, within the category of “The most frequent errors”, AssertionError and NameError exhibit distinctly different trends: NameError is relatively evenly distributed while the frequency of AssertionError significantly decreases as the model’s accuracy improves. To verify this, we calculated the Pearson correlation coefficients between the number of these two types of errors and the corresponding model’s accuracy. The results reveal a strong negative correlation between AssertionError occurrences and model accuracy, with a Pearson correlation coefficient of -0.84 and a highly significant p-value of $1.7\times 10^{-4}$. This suggests that as model accuracy improves, the frequency of AssertionErrors decreases substantially. In contrast, NameError shows a weaker negative correlation with model accuracy, having a correlation coefficient of -0.41 and a non-significant p-value of 0.14. This indicates that while there is a slight tendency for NameErrors to decrease as accuracy improves, the relationship is not statistically significant and the trend is less pronounced compared to AssertionErrors. For ”Occasionally occurring errors,” we also conducted a similar correlation analysis and found that SyntaxError and IndexError do not show a significant correlation, while the Pearson correlation coefficients for TypeError and ValueError are around -0.7, indicating a strong negative correlation. Since errors in the “Rarely occurring errors” category hardly ever occur, we could not conduct more in-depth statistical analysis on them.
From the model’s perspective, we found that only GPT-3.5-turbo and Phi-2-2.7B encountered ModuleNotFoundError, indicating that they not only called modules from the Python standard library but might also have used modules from non-standard libraries to aid in problem-solving. Additionally, GPT-3.5-turbo had the fewest number of AssertionErrors and the highest number of NameErrors, mainly because it often calls modules directly without adding import statements. This point is a significant consideration in our proposed solution, indicating that our approach is likely to be very effective in this model. Moreover, we also found that GPT-3.5-turbo and Phind-CodeLlama-V2-34B rarely experience SyntaxError, possibly because they are more proficient in mastering Python syntax and generally do not generate unnecessary functions, thus avoiding SyntaxErrors caused by incomplete functions. In contrast, Phi-1.5-1.3B and Phi-2-2.7B exhibited a higher frequency of SyntaxError occurrences. This could be attributed to their tendency to generate more complex or convoluted code structures, leading to a greater likelihood of syntax-related issues such as incomplete or incorrect function definitions.
III-D4 The relationship between failed samples and errors
A single piece of code may contain multiple errors. However, the Analysis Process described in III-D is currently limited to extracting only the first error information from the code. To address this limitation, we propose a manual analysis process, illustrated in Fig. 8. This process aims to identify all errors in the code through an iterative “fix-test” cycle. By applying this process to code generated by LLMs, we confirmed that a single code snippet can contain multiple errors, establishing a one-to-many relationship between failed code snippets and errors. For example, the prompt from the HumanEval dataset (task_id 133), which serves as the comment in Listing LABEL:list_ft1, specifies the task for the LLM: squaring each number in a given list of digits and rounding each element to the upper integer (ceiling) first. The original code generated by the model, shown in Listing LABEL:list_ft1, triggered a SyntaxError due to an unclosed “[”. Based on the error type and the information provided, we manually corrected the error, resulting in the code shown in Listing LABEL:list_ft2. Subsequent testing revealed that this corrected code still failed to pass all tests. The error occurred because a generator expression was passed into the int() function, which expects a string, bytes, or a real number, leading to a TypeError. Consequently, we performed a second manual error correction, resulting in the code shown in Listing LABEL:list_ft3. After this final correction, the code successfully passed all test cases.
Figure 8: Schematic Diagram of the “Fix-Test” Process.
{mdframed}
[nobreak=true] Answer to RQ2: We collected a total of 12,837 errors, which can be classified into 14 categories according to Python’s built-in exception types. Among these, AssertionError emerged as the most common error type. The causes of these errors can be attributed to 19222The “Error Cause” column in Table III has 20 rows, but both ValueError and NotImplementError can be caused by the same cause, “Intentional Raise”. Therefore, there are only 19 distinct error causes in total. different causes, such as incorrect function implementation, syntax mistakes, logical errors, and failure to handle edge cases.
III-E Which Errors Can be Directly Fixed? (RQ3)
III-E1 Motivation and Method
This subsection describes the process of manually analyzing errors from LLM-generated code to determine which types are suitable for automated fixing. The approach involves sampling from the errors obtained from the test on the HumanEval dataset, analyzing these samples, and then establishing criteria for automatically fixable errors.
We tested 14 LLMs, running each model ten times on the HumanEval dataset. This process resulted in a total of 12,837 errors. For each error type, we employed a stratified sampling approach to ensure comprehensive representation. For categories with fewer than 100 samples, we included all available examples. For categories with 100 or more samples, we sampled 30% of the total, with a minimum of 100 and a maximum of 200 samples per category. This adaptive sampling strategy allowed us to maintain a balance between thoroughness and manageability across varied error frequencies.
After analyzing these sampled errors, we engaged in extensive discussions between two authors and solicited a wide range of opinions. We unanimously established the following criteria for errors that can be automatically fixed:
  1. $\bullet$
    Consistency and Predictability. The error exhibits high consistency and predictability, manifesting similarly across different code instances, and can be addressed using general rules or patterns. This contrasts with errors that are highly dependent on specific code logic and require an understanding of particular code context for resolution.
  2. $\bullet$
    Locality. The error is confined to a specific part of the code and requires only local context information for fixing. This is in contrast to errors that may involve interactions across an entire function or between multiple functions, necessitating broader context or external knowledge, such as understanding the overall problem logic.
  3. $\bullet$
    Complexity of Fix. The error can be resolved through simple rules or pattern matching. This is opposed to errors requiring complex reasoning or multi-step modifications for correction.
TABLE IV: Analysis of error causes for automatic fixability based on criteria mentioned in subsection III-E
      Error Cause       Consistency and Predictability       Locality       Low Complexity       Fixable?
      Test Case Failure       $\times$       $\times$       $\times$       $\times$
      Empty Function       ✓       ✓       $\times$       $\times$
      Misremembered Name       $\times$       $\times$       $\times$       $\times$
      Missing Content       ✓       ✓       $\times$       $\times$
      Missing Import       ✓       ✓       ✓       ✓
      Unbalanced Delimiters       $\times$       ✓       $\times$       $\times$
      Function Overflow       ✓       ✓       ✓       ✓
      Empty Sequence       $\times$       ✓       $\times$       $\times$
      Intentional Raise       $\times$       ✓       $\times$       $\times$
      Inappropriate Argument       $\times$       ✓       $\times$       $\times$
      Out of Bounds       $\times$       ✓       $\times$       $\times$
      Incompatible Operation       $\times$       $\times$       $\times$       $\times$
      Non-existent Attribute       $\times$       ✓       $\times$       $\times$
      Execution Timeout       $\times$       $\times$       $\times$       $\times$
      Inconsistent Indentation       ✓       ✓       ✓       ✓
      Missing Module       $\times$       ✓       ✓       $\times$
      Non-existent Key       $\times$       $\times$       $\times$       $\times$
      Unassigned Variable       $\times$       $\times$       $\times$       $\times$
      Infinite Recursion       $\times$       $\times$       $\times$       $\times$
III-E2 Error Analysis Results
Based on the above criteria, we analyzed the fixability of each error type and its corresponding reasons. The results are summarized in Table IV. A detailed explanation of this table is as follows:
Three types of errors - Missing Import, Function Overflow, and Inconsistent Indentation - demonstrate high consistency and predictability. For instance, they consistently manifest as missing imports for specific modules, incomplete functions at the end of the code, or indentation issues. These errors also exhibit high locality, such as incomplete functions appearing only at the end of the code. Moreover, the complexity of fixing these issues is relatively low. For example, missing import statements can be added at the beginning of the code, incomplete functions at the end can be removed, and indentation can be standardized uniformly. Therefore, these three types of errors can be automatically fixed.
While Empty Function and Missing Content errors show high consistency, predictability, and locality, they fail from the absence of crucial code. Since we cannot predict how a function should implement a given functionality, we cannot automatically generate the missing content, resulting in extremely high complexity for fixing these issues. Consequently, these two types of errors cannot be automatically fixed.
Regarding Missing Module errors, although they exhibit locality and can be fixed by installing the corresponding module version, we cannot predict which non-standard library modules and versions the model will use. Thus, this type of error lacks predictability and cannot be automatically fixed.
Six types of errors - Unbalanced Delimiters, Empty Sequence, Intentional Raise, Inappropriate Argument, Out of Bounds, and Non-existent Attribute - demonstrate high locality as they typically involve only a small portion of the code. However, they lack consistency and predictability due to varying causes in different tasks. Fixing these errors requires understanding the code structure and intent, which is nearly impossible for automatic fix, resulting in high complexity. Therefore, they cannot be automatically fixed.
Seven types of errors - Test Case Failure, Misremembered Name, Incompatible Operation, Execution Timeout, Non-existent Key, Unassigned Variable, and Infinite Recursion - usually involve specific logic and functionality of the code. They require understanding the specific requirements of the problem and the expected output of test cases, and cannot be resolved with simple rules. These errors do not meet the criteria of consistency and predictability, locality, and low complexity, and thus cannot be automatically fixed.
{mdframed}
[nobreak=true] Answer to RQ3: Errors amenable to automated fixing should meet three crucial criteria: high consistency and predictability, high locality, and low complexity of fix. Based on these criteria, we identified three types of errors suitable for automated fixing including Missing Import, Function Overflow, and Inconsistent Indentation.
IV Methodology
RQ3 indicated that errors caused by three reasons can be directly and automatically fixed. This section details the proposed fixing method named LlmFix for such errors, which is further evaluated in Section V.
IV-A Overall Framework of LlmFix
Generally, LlmFix addresses errors in three distinct steps: Code Filtering (Step-1), Code Truncation (Step-2) and Importing Missing Modules (Step-3). Code Filtering focuses on rectifying Inconsistent Indentation, Code Truncation focuses on rectifying Function Overflow, while Importing Missing Modules concentrates on resolving Missing Import issues. Implementation details are described below. Generally, we have integrated the various fix methods mentioned previously to construct a complete fix process LlmFix, as shown in Fig. 9. To enhance understanding of this process, we have also created an example that demonstrates the specific function of each step within this fix process, as detailed in Fig. 10.
Figure 9: Flowchart of the method LlmFix.
Figure 10: An example of applying LlmFix. This example was crafted by us to help readers understand better. In practice, not all the fix steps are used within the same code segment.
IV-B Step-1: Code Filtering
In addition to the __main__ block, the code generated by the model may also contain input statements and assert statements. The former can cause the test program to get stuck waiting for user input, while the latter can lead to unnecessary errors since the test cases generated by the model itself might be incorrect. Therefore, we need to filter these out. Moreover, as shown in Listing LABEL:list_norm1 in Appendix, an unequal number of spaces used for indentation could cause an IndentationError. To avoid this issue, we replace all space-based indentations with tabs.
IV-C Step-2: Code Truncation
We have observed that many LLMs tend to generate multiple functions until reaching the set length limit, rather than just generating the necessary function. In this scenario, some functions reach the generation length limit before they are fully implemented, leaving an incomplete function structure that causes a SyntaxError.
To address this, two common solutions are employed: one is to retain only the code of the first function; the other is to remove the last incomplete function. However, neither method achieves ideal results. For example, as shown in Listing LABEL:list_appr3 in Appendix, if only the first function’s code is retained, the entry function prime_fib() would be removed, leading to NameError; similarly, in Listing A, functions encode_vigenere() and decode_vigenere() cause a SyntaxError due to unenclosed brackets. These functions are not called by the entry function encode_cyclic(), yet they cause the entire code segment containing the correct code to fail compilation.
This shows that using any single method alone cannot completely avoid above issues. Therefore, we propose a new truncation strategy that combines the ideas of both methods. The process starts by compiling the untruncated code directly; if compilation fails, the last incomplete function is removed before attempting compilation again. If compilation still fails, this indicates a situation similar to that shown in Listing A might be occurring, hence further removal of functions at the end of the code is performed for compiling. This process is repeated until only one function remains or the code compiles successfully, at which point the outputted code is considered the optimal truncation result. This truncation method not only avoids issues caused by deleting too many functions but also prevents Function Overflow from causing the entire segment of code to fail to compile, thus more accurately demonstrating the model’s performance.
IV-D Step-3: Importing Missing Modules
We have noticed that the model often calls certain modules directly without adding an import, and most of these modules are from the Python standard library. Based on this observation, we have built a mini-database that contains the names of modules frequently called by the model, as well as the names of variables and functions within those modules. When a NameError occurs, we refer to the error message indicating name ’XXX’ is not defined. (for example, name ’math’ is not defined.), and we look up this name in the previously constructed mini-database. If it is included, we add its import statement.
For example, for the sample shown in Listing LABEL:list_import1, the model directly called the hashlib module from the Python standard library without importing it, which triggered a NameError and returned the error message name ’hashlib’ is not defined. In this case, our program will detect that the NameError was caused by hashlib, which is a name that exists in the Python standard library, and therefore, an import statement for hashlib will be added at the beginning of the generated code to prevent this issue, as shown in the modified code in Listing LABEL:list_import2 in Appendix. After correctly importing the hashlib module, since the original code’s logic was correct, the modified code successfully passed all test cases.
V Evaluation
V-A Research Question
RQ4: How effective is LlmFix? This RQ conducted comparative tests on the 14 LLMs tested in the empirical study, using both the HumanEval and MBPP datasets, under scenarios with and without the use of LlmFix. We also statistically analyzed the improvement LlmFix brings to different LLMs and the proportion of different error types it successfully fixes.
V-B Datasets
We apply the LlmFix to run ten additional times on each of the 14 models selected in the empirical study using both the HumanEval and MBPP datasets’ test sets. This approach aims to demonstrate that the method is not only effective on HumanEval but also equally effective on the previously unstudied MBPP dataset, thus showcasing a certain degree of generalization capability. The MBPP (Mostly Basic Python Problems) dataset is a comprehensive collection of 974 Python programming tasks designed to evaluate code generation and program synthesis capabilities [60]. Each task in the dataset is accompanied by an average of 3 test cases, ensuring the functional correctness of the generated code. Prominent language models such as GPT-3 [26], Codex [13], and PaLM [61] have utilized MBPP for assessing their code generation proficiency.
As illustrated in Fig. 11, the MBPP dataset comprises six essential components: 1) task_id, a unique identifier assigned to each task, facilitating easy reference and organization of the code generation challenges. 2) text, a paragraph of text that provides generation requirements for guiding LLMs. 3) code, the standard solution for each task. 4) test_list, a testing list including several test cases. 5) test_setup_code, a piece of code that runs before executing the tests, ensuring that the tests are executed in the correct context. 6) challenge_test_list, a set of additional, more challenging test cases used to deeply evaluate the robustness and correctness of the generated code.
Figure 11: A schematic diagram of the first sample in MBPP test sets.
V-C Effectiveness of LlmFix in Error Fixing
Table V lists the specific numerical characteristics of the ten tests conducted with LlmFix across 14 LLMs in the HumanEval dataset, including minimum, maximum, mean, and standard deviation. Comparing with Table I, we found that with LlmFix, the average test results of each model improved by about 9.5%, with a maximum improvement of 16.46% percentage points. In terms of a total number of errors, LlmFix addressed 17.07% of all errors in HumanEval (reducing errors from 12,837 to 10,646). Additionally, out of the ten models that publicly reported their performance, six of them surpassed the reported SOTA performance using our fixing process. We performed a Wilcoxon signed-rank test on the mean test results before and after fixing, obtaining a p-value of $1.22\times 10^{-4}$, which is much smaller than the significance level of 0.05. This result indicates that our fixing process significantly improved the performance of the models. Furthermore, the standard deviation (STD) of all models is much lower than their average value (less than 10% of the average), indicating very low data variability.
TABLE V: Accuracy of ten tests (with LlmFix) of LLMs on HumanEval dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values.
Model (With LlmFix) Min Max Mean STD
CodeLlama-Python-7B 37.80 43.29 39.88 1.85
CodeLlama-Python-13B 42.68 46.95 44.51 1.32
CodeLlama-Python-34B 52.44 56.71 54.63 1.63
WizardCoder-Python-V1.0-7B 50.00 54.88 51.52 1.44
WizardCoder-Python-V1.0-13B 54.88 56.71 55.91 0.71
WizardCoder-Python-V1.0-34B 66.46 69.51 67.44 0.96
Phind-CodeLlama-V2-34B 65.85 74.39 71.40 2.34
Phi-1-1.3B 50.00 52.44 51.04 0.71
Phi-1.5-1.3B 38.41 43.90 40.43 1.80
Phi-2-2.7B 49.39 54.27 52.01 1.55
GPT-3.5-turbo 70.12 75.00 72.62 1.51
CodeTulu-2-7B 40.24 48.78 43.84 2.10
CodeTulu-2-13B 45.73 55.49 50.55 2.59
CodeTulu-2-34B 53.05 57.32 55.06 1.47
We have identified that among all 19 error causes, only three causes are amenable to automatic fixing: Missing Imports, Function Overflow, and Inconsistent Indentation. Among them, Missing Import is one of the causes leading to NameError, Function Overflow is one of the causes leading to SyntaxError, and Inconsistent Indentation is one of the causes leading to IndentationError. Given that compiling statistics based on error types is more convenient, and considering the low frequency of IndentationErrors and their similarity to SyntaxErrors in terms of root causes [62], we have incorporated the IndentationError statistics into the SyntaxError category here. Subsequently, we compared the frequency of NameErrors and SyntaxErrors before and after applying LlmFix. The results of this comparison are illustrated in Fig. 12. We found that our proposed fixing process fixed an average of 90.26% of NameErrors and 99.74% of SyntaxErrors.
Figure 12: The average sample number of all NameErrors and SyntaxErrors (including IndentationErrors) per round of tests using the HumanEval dataset, without or with LlmFix.
Table VI shows the specific numerical characteristics of the ten tests conducted across 14 LLMs in the MBPP dataset, comparing scenarios with and without LlmFix, including minimum, maximum, mean, and standard deviation. We found that with LlmFix, the average test results of each model improved by about 5.4%, with a maximum improvement of 8.48%, on average. We performed a Wilcoxon signed-rank test on the mean test results before and after fixing, obtaining a p-value of $1.22\times 10^{-4}$, which is much smaller than the significance level of 0.05. This result indicates that LlmFix significantly improved the performance of the models. Furthermore, the standard deviation (STD) of all models is much lower than their average value (less than 10% of the average), indicating very low data variability.
Model (Without LlmFix) Min Max Mean STD
CodeLlama-Python-7B 40.60 43.40 42.12 0.94
CodeLlama-Python-13B 45.60 48.40 46.86 0.92
CodeLlama-Python-34B 47.80 51.00 49.82 1.03
WizardCoder-Python-V1.0-7B 48.60 51.40 50.08 0.92
WizardCoder-Python-V1.0-13B 49.80 52.20 50.76 0.79
WizardCoder-Python-V1.0-34B 55.60 57.40 56.68 0.53
Phind-CodeLlama-V2-34B 56.20 58.80 57.16 0.91
Phi-1-1.3B 46.20 47.80 47.02 0.55
Phi-1.5-1.3B 34.00 38.60 36.40 1.46
Phi-2-2.7B 41.80 44.00 42.70 0.73
GPT-3.5-turbo 63.20 66.40 64.64 1.02
CodeTulu-2-7B 43.60 45.40 44.50 0.63
CodeTulu-2-13B 47.40 50.20 49.02 0.80
CodeTulu-2-34B 53.00 56.40 54.38 0.92
Model (With LlmFix) Min Max Mean STD
CodeLlama-Python-7B 47.20 49.80 48.04 0.82
CodeLlama-Python-13B 51.00 54.20 52.46 1.01
CodeLlama-Python-34B 56.20 58.20 57.24 0.74
WizardCoder-Python-V1.0-7B 54.00 55.40 54.42 0.47
WizardCoder-Python-V1.0-13B 53.00 54.80 54.00 0.63
WizardCoder-Python-V1.0-34B 60.00 62.60 61.12 0.83
Phind-CodeLlama-V2-34B 60.20 63.20 61.44 0.85
Phi-1-1.3B 47.60 50.80 49.32 0.95
Phi-1.5-1.3B 42.40 45.00 43.86 0.85
Phi-2-2.7B 50.00 53.00 51.18 0.91
GPT-3.5-turbo 70.40 73.40 72.18 0.87
CodeTulu-2-7B 46.80 51.20 49.36 1.43
CodeTulu-2-13B 51.40 55.40 53.52 1.21
CodeTulu-2-34B 57.80 60.60 59.30 0.86
TABLE VI: Accuracy of ten tests(without/with LlmFix) of LLMs on MBPP dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values.
Similar to the research for HumanEval, we plotted the bar chart shown in Fig. 13 to visually present the average number of samples per round of testing for all NameErrors and SyntaxErrors (including IndentationError) with LlmFix and without LlmFix. We found that LlmFix fixed an average of 91.47% of NameErrors and 100% of SyntaxErrors.
Figure 13: The average sample number of all NameErrors and SyntaxErrors (including IndentationErrors) per round of tests using the MBPP dataset, without or with LlmFix.
{mdframed}
[nobreak=true] Answer to RQ4: Our proposed method, LlmFix, demonstrates the capability to resolve nearly all errors that are amenable to automated fixing in both the HumanEval and MBPP datasets. LlmFix can provide 9.5% and 5.4% average accuracy improvements for 14 LLMs on HumanEval and MBPP, respectively.
VI Discussion
VI-A Implication
Based on our experimental results and discussion, this section provides implications for model developers and future research.
For model developers, when introducing a new LLM, it is beneficial to concurrently develop a customized version of LlmFix, thereby further enhancing the model’s practical performance. We developed a general-purpose LlmFix based on the experimental results of 14 LLMs on the HumanEval dataset, as discussed in Section V. This version has demonstrated excellent performance on both the HumanEval and MBPP datasets, thus serving as a benchmark tool applicable to various models and datasets. However, to achieve better results, model developers can fine-tune the baseline LlmFix specifically for their models, potentially leading to more significant performance improvements for those particular models.
For future research directions, LlmFix can serve as a novel metric for assessing code quality. We can posit that code passing tests without requiring LlmFix processing is of higher quality than code needing LlmFix treatment to pass tests. As mentioned in subsection IV-A, LlmFix comprises three steps, though not all code requires all three steps as illustrated in Fig. 10. Building on this observation, we can further infer that when $m<n$, code passing tests after only $m$ steps of LlmFix processing is of higher quality than code requiring $n$ steps to pass tests. This provides a new perspective for a more nuanced evaluation of code quality.
VI-B Threats to Validity
There are some potential threats affecting the validity of our experimental results and conclusions.
Limited Dataset. In our research, we evaluated the code generation accuracy of 14 selected models. Due to our limited computation resources, we focused our testing on two widely-used datasets: HumanEval and MBPP. While these datasets are well-established in the field, the narrow scope of our data selection raises questions about the generalizability of our experimental results and findings. The limited dataset diversity leaves room for uncertainty regarding how our conclusions might apply to other, potentially more diverse, or specialized, coding datasets. This limitation underscores the need for future studies to incorporate a broader range of datasets to validate and potentially extend our findings.
Limited Testing. Our study focused on testing the Pass@1 results of 14 LLMs on the HumanEval and MBPP datasets, leading to the development and validation of LlmFix. However, due to constraints in cost and time, we were unable to conduct tests for Pass@10 or even Pass@100. Extending our experiments to include these higher sampling numbers would undoubtedly make our study more comprehensive, thereby strengthening the persuasiveness and reliability of our conclusions. This highlights a potential area for improvement in future research.
VII Conclusion
In this study, we conducted ten rounds of tests on 14 LLMs (such as codeLlama-Python, WizardCoder-Python, and GPT-3.5-turbo) using the HumanEval dataset. Through manual analysis of the test results, we found that these LLMs achieved an average of 84.07% of their reported performance, demonstrating high reproducibility. We also investigated the relationship between Pass@1 results, model inference time, and model parameter size. Our analysis revealed a positive correlation between Pass@1 results and model parameter size, while no significant correlation was observed between inference time and parameter size. Subsequently, we performed an in-depth analysis of errors in the test results, extracting and categorizing 12,837 errors into 14 types. Through in-depth manual analysis, we ultimately identified 19 specific causes leading to these errors. Based on established evaluation criteria, we assessed whether each error cause could be automatically fixed, leading to our development of LlmFix, a method designed to address code generation errors. Evaluation of LlmFix on both HumanEval and MBPP datasets showed that it effectively resolved almost all automatically fixable errors. LlmFix can improve the performance of 14 LLMs by an average of 9.5% and 5.4% on HumanEval and MBPP, respectively.
Appendix A Code Examples Repository
[…]
2def string_to_md5(text):
3 if text == ’’:
4 return None
5 else:
6 return hashlib.md5(text.encode(’utf-8’))
7 .hexdigest()
[…]
2import hashlib
3
4def string_to_md5(text):
5 if text == ’’:
6 return None
7 else:
8 return hashlib.md5(text.encode(’utf-8’))
9 .hexdigest()
1def fruit_distribution(s,n):
2 apples, oranges = s.split(" ")
3 apples = int(apples.split("apples")[0])
4 oranges = int(oranges.split("oranges")[0])
5 return n - apples - oranges
[…]
2def play_hand(hand: str, word_list: List[str])->None:
3
4 score = 0
5
6 while calculate_handlen(hand) > 0:
7
8 display_hand(hand)
9 user_word = input(’Enter word, or a "." to indicate that you are finished: ’)
10
11 if user_word == ’.’:
12 break
13
14 if is_valid_word(user_word, hand, word_list):
15 score += get_word_score(user_word,
16 calculate_handlen(hand))
17 hand = update_hand(hand, user_word)
[…]
2def vigenere_cipher(s: str, k: str):
3 alphabet = abc
[…]
2def is_prime(num):
3 if num < 2:
4 return False
5 for i in range(2, int(num ** 0.5) + 1):
6 if num % i == 0:
7 return False
8 return True
9
10def prime_fib(n: int):
11 if n == 1:
12 return 2
13 elif n == 2:
14 return 3
15 fib_list = [2, 3]
16 i = 2
17 while len(fib_list) < n:
18 next_fib = fib_list[i-1] + fib_list[i-2]
19 if is_prime(next_fib):
20 fib_list.append(next_fib)
21 i += 1
22 return fib_list[-1]
1def numerical_letter_grade(grades):
2 pass
1def intersection(interval1, interval2):
2 if interval1[1] < interval2[0] or interval2[1] <
3 interval1[0]:
4 return "NO"
5 else:
6 start = max(interval1[0], interval2[0])
7 end = min(interval1[1], interval2[1])
8 length = end - start
9 if isPrime(length):
10 return "YES"
11 else:
12 return "NO"
13
14def is_prime(num):
15 if num == 1:
16 return False
17 else:
18 for i in range(2, num):
19 if num % i == 0:
20 return False
21 return True
1def encode_cyclic(s: str):
2 # split string to groups. Each of length 3.
3 groups = [s[(3 * i):min((3 * i + 3), len(s))] for
4 i in range((len(s) + 2) // 3)]
5 # cycle elements in each group. Unless group has
6 # fewer elements than 3.
7 groups = [(group[1:] + group[0]) if len(group) ==
8 3 else group for group in groups]
9 return "".join(groups)
10
11
12def decode_cyclic(s: str):
13 # split string to groups. Each of length 3.
14 groups = [s[(3 * i):min((3 * i + 3), len(s))] for
15 i in range((len(s) + 2) // 3)]
16 # cycle elements in each group. Unless group has
17 # fewer elements than 3.
18 groups = [(group[1:] + group[0]) if len(group) ==
19 3 else group for group in groups]
20 return "".join(groups)
21
22def encode_caesar(s: str, k: int):
23 return "".join([chr((ord(c) - 97 + k) % 26 + 97)
24 for c in s])
25
26def decode_caesar(s: str, k: int):
27 return "".join([chr((ord(c) - 97 - k) % 26 + 97)
28 for c in s])
29
30def encode_vigenere(s: str, k: str):
31 # convert key to lowercase
32 k = k.lower()
33 # repeat key until it is as long as the message
34 k = k * ((len(s) + len(k) - 1) // len(k))
35 # apply vigenere cipher to each character in s
36 return "".join([chr(((ord(c) - 97 + ord(k[i]) -
37 97) % 26 + 97) for c in s])
38
39def decode_vigenere(s: str, k: str):
40 # convert key to lowercase
41 k = k.lower()
42 # repeat key until it is as long as the message
43 k = k * ((len(s) + len(k) - 1) // len(k))
44 # apply vigenere cipher to each character in s
45 return "".join([chr(((ord(c) - 97 - ord(k[i]) +
46 26) % 26 + 97) for c in s])
47
48def encode_substitution(s: str, d: dict):
49 return "".join([d[c] for c in s])
50
51def decode_substitution(s: str, d: dict):
52 return "".join([c for c in s if c in d])
1def how_many_times(string: str, substring: str)->int:
2 return len(string) - len(string.replace(substring,
3 ’’, len(string)))
1def same_chars(s0: str, s1: str):
1def string_to_md5(text):
2 if text == ’’:
3 return None
4 else:
5 return hashlib.md5(text.encode(’utf-8’))
6 .hexdigest()
1def common(l1: list, l2: list):
2 return sorted(list(set(l1) & set(l2))
1def add(x: int, y: int):
2 return x + y
3
4...
5
6def vigenere_cipher(s: str, k: str):
7 alphabet = abc
1def strange_sort_list(lst):
2 if not lst:
3 return []
4 min_val = min(lst)
5 max_val = max(lst)
6 result = [min_val]
7 lst.remove(min_val)
8 lst.remove(max_val)
9 result.append(max_val)
10 return result + strange_sort_list(lst)
1import math
2
3def poly(xs: list, x: float):
4 return sum([coeff * math.pow(x, i) for i, coeff in enumerate(xs)])
5
6def find_zero(xs: list):
7 if len(xs) % 2!= 0:
8 raise ValueError("xs must have even number of coefficients")
9 if xs[0]!= 0:
10 raise ValueError("largest non zero coefficient must be 0")
11 x = 0
12 while poly(xs, x)!= 0:
13 x += 1
14 return x
1def check_if_last_char_is_a_letter(txt):
2 if txt[-1].isalpha() and not txt[-2].isalnum():
3 return True
4 else:
5 return False
1def sort_array(array):
2 if not array:
3 return []
4 if len(array) == 1:
5 return array
6 if sum(array[0], array[-1]) % 2 == 0:
7 return sorted(array)
8 else:
9 return sorted(array, reverse=True)
1def check_dict_case(dict):
2 if len(dict) == 0:
3 return False
4 else:
5 return all(k.islower() for k in dict.keys()) or all(k.isupper() for k in dict.keys())
1def minPath(grid, k):
2 n = len(grid)
3 dp = [[float(’inf’)] * n for _ in range(n)]
4 dp[0][0] = 0
5 for i in range(n):
6 for j in range(n):
7 if i > 0:
8 dp[i][j] = min(dp[i][j], dp[i-1][j]
9 + grid[i][j])
10 if j > 0:
11 dp[i][j] = min(dp[i][j], dp[i][j-1]
12 + grid[i][j])
13 path = []
14 i = n-1
15 j = n-1
16 while i >= 0 and j >= 0:
17 path.append(grid[i][j])
18 if i > 0 and dp[i-1][j] == dp[i][j]
19 + grid[i][j]:
20 i -= 1
21 elif j > 0 and dp[i][j-1] == dp[i][j]
22 + grid[i][j]:
23 j -= 1
24 path.reverse()
25 return path[:k]
[…]
2...
3
4def play_hand(hand, word_list):
5 total_score = 0
6 while calculate_handlen(hand) > 0:
7 display_hand(hand)
8 word = input("Enter word, or a ’.’ to indicate that you are finished: ")
9 if word == ".":
10 break
11 elif is_valid_word(word, hand, word_list):
12 score = get_word_score(word,
13 calculate_handlen(hand))
14 total_score += score
15
16...
[…]
2import roman
3
4def int_to_mini_roman(number):
5 return roman.toRoman(number).lower()
[…]
2from typing import List
3
4
5def parse_music(music_string: str) -> List[int]:
6 notes = {’o’: 4, o|’: 2, ’.|’: 1}
7 music = music_string.split(’ ’)
8 return [notes[note] for note in music]
[…]
2def is_palindrome(string: str) -> bool:
3 return string == string[::-1]
4
5def make_palindrome(string: str) -> str:
6 if not string:
7 return string
8
9 # Find the longest palindromic suffix of the
10 # string
11 i = 1
12 while i <= len(string) // 2:
13 prefix = string[:-i]
14 suffix = string[-i:]
15 if is_palindrome(suffix):
16 break
17 i += 1
18
19 # Reverse the prefix
20 prefix = prefix[::-1]
21
22 return prefix + string
[…]
2def special_factorial(n):
3 if n == 1:
4 return 1
5 else:
6 return n * special_factorial(n-1)
7 * special_factorial(n-2)
[…]
2from typing import List
3
4def separate_paren_groups(paren_string: str) ->
5 List[str]:
6 # TODO: implement this function
7 raise NotImplementedError
[…]
2def sum_squares(lst):
3 """
4 You are given a list of numbers.
5 You need to return the sum of squared numbers in
6 the given list, round each element in the list
7 to the upper int(Ceiling) first.
8 Examples:
9 For lst = [1,2,3] the output should be 14
10 For lst = [1,4,9] the output should be 98
11 For lst = [1,3,5,7] the output should be 84
12 For lst = [1.4,4.2,0] the output should be 29
13 For lst = [-2.4,1,1] the output should be 6
14 """
15 lst = [int(math.ceil(num**2) for num in lst)
16 return sum(lst)
[…]
2def sum_squares(lst):
3 lst = [int(math.ceil(num**2) for num in lst)]
4 return sum(lst)
[…]
2import math
3
4def sum_squares(lst):
5 lst = [math.ceil(num)**2 for num in lst]
6 return sum(lst)